OpenStudy (anonymous):

Come here : @Jack162

OpenStudy (anonymous):

Now proceed from beginning of this concept..

OpenStudy (anonymous):

Ok for any prime number p, define $$v_p(n)$$ To be the largest prime that divides n do you see why $$v_p(ab)=v_p(a)+v_p(b)$$

OpenStudy (anonymous):

For natural numbers a and b?

OpenStudy (anonymous):

Yes I got it..

OpenStudy (anonymous):

Ok then we know that $$v_p(m!)=v_p(1)+v_p(2)+v_p(3)+...v_p(m)$$

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

Written more concisely $$v_p(m!)=\sum_{n=1}^m v_p(n)$$

OpenStudy (anonymous):

Now suppose p does not divide n

OpenStudy (anonymous):

Getting..

OpenStudy (anonymous):

What will v_p(n) be?

OpenStudy (anonymous):

Don't know..

OpenStudy (anonymous):

$$n!=1\times 2\times 3\times 4\times 5\times 6.... \times n$$

OpenStudy (anonymous):

How many multiples of p are there in n!?

OpenStudy (anonymous):

We have $$p,2p,3p,...p[\frac{n}{p}]$$

OpenStudy (anonymous):

Where [.] denotes the floor function

OpenStudy (anonymous):

Okay..

OpenStudy (anonymous):

do you understand why? the closest multiple of p to n is p[n/p]

OpenStudy (anonymous):

Now looking at the rest of the factors in n!, they are not divisible by p so we can exclude them as they have no factor of p.

OpenStudy (anonymous):

Thus we need only count the factors of p in$$p*2p*3p*4p*5p*6p*7p*8p*...p[\frac{n}{p}]$$

OpenStudy (anonymous):

Okay, getting..

OpenStudy (anonymous):

Factoring out p we see that thats equivilent to $$p^{[\frac{n}{p}]}1*2*3*4*5*...*[\frac{n}{p}]$$

OpenStudy (anonymous):

Now lets look at the multiples of p in $$1*2*3*4*...*[\frac{n}{p}]$$

OpenStudy (anonymous):

OKay..

OpenStudy (anonymous):

They are $$p,2p,3p,4p,...p[\frac{n}{p^2}]$$

OpenStudy (anonymous):

But sense everything else in the product isn't divisible by p, they each contribute no factors so we can exclude them and only keep the multiples of p.

OpenStudy (anonymous):

So again you will take P raise to the power n/p^2 common??

OpenStudy (anonymous):

yes, and we then do it with $$[n/p^3],[n/p^4],[n/p^5]$$ So in the end we just need to count how many factors of p, $$p^{[n/p]+[n/p^2]+[n/p^3]...}$$ Has, Which is clearly $$[\frac{n}{p}]+[\frac{n}{p^2}]+[\frac{n}{p^3}]+...$$

OpenStudy (anonymous):

So we know n! has that many factors of p

OpenStudy (anonymous):

Oh, this look similar to formula given by Directrix.

OpenStudy (anonymous):

Yes you can use similar reasoning to find the greatest dividing exponent in many other products

OpenStudy (anonymous):

So, how to proceed further with this to solve 40!^{40!)

OpenStudy (anonymous):

Ok well look if n has say k factors p how many factors will n^a have? it will a times as many factors sense were raising it to the ath power

OpenStudy (anonymous):

The same reasoning is used here

OpenStudy (anonymous):

Sense were raising 40! to 40!, it will have 40! times the number of factors of 10 40! has

OpenStudy (anonymous):

Can you explain it more ??

OpenStudy (anonymous):

Ok heres an example, how many factors of 2 does $$(5*2*3*7)^{100}$$ have?

OpenStudy (anonymous):

Or how many factors of 5 does, $$(7*5^4*3)^{10}$$ have?

OpenStudy (anonymous):

Sense were raising it to a power, it will that times the number of factors on the inside of the expression.

OpenStudy (anonymous):

So say we wrote $$40!=10^a*p_1*p_2*p_3$$ How many factors of 10 will $$40!^{40!}$$ have?

OpenStudy (anonymous):

Do you see why its 40!a

OpenStudy (anonymous):

Assuming none of the primes on the right contribute a factor of 10

OpenStudy (anonymous):

Not getting.. Wait, you said the power is the number of factors inside the expression, right??

OpenStudy (anonymous):

We want to find the number of trailing zeros in 40!^40!, this is the same thing as counting how many factors of 10 it has

OpenStudy (anonymous):

Yes.. This I understand..

OpenStudy (anonymous):

Do you see why?

OpenStudy (anonymous):

Ok so we need to find out how many factors of 10 it has

OpenStudy (anonymous):

Yes,,

OpenStudy (anonymous):

Lets say we know how many factors of 10, 40! has call it a, then we can write $$40!=10^{a}*p_1*p_2*p_3*p_4*...$$

OpenStudy (anonymous):

Yes we can..

OpenStudy (anonymous):

How many factors of 10 does $$40^{40!}=10^{a*40!}*p_1^{40!}*p_2^{40!}...$$ have now?

OpenStudy (anonymous):

It has $$40!\times a$$ Factors of 10

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

So we just need to find a, what did we say a was again?

OpenStudy (anonymous):

Zeros in 40!

OpenStudy (anonymous):

Ok so 40!^{40!} has $$40!\times (\text{# of zeros in 40!)}$$ Factors of 10

OpenStudy (anonymous):

I mean factors of 10..

OpenStudy (anonymous):

same thing as we said before

OpenStudy (anonymous):

Okay, 9 factors of 10 are therer. ie 10^9

OpenStudy (anonymous):

So, \(\color{green}{9 \times 40 !}\)

OpenStudy (anonymous):

So we know it has $$40!^{40!}$$ has 9*40! trailing zeros

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

This is Arithmetic at its best.. Wow.. Can we go with now highest power of 3 in 80! ??

OpenStudy (anonymous):

Yes use the formula we proved a while back here, that says how many factors of a prime p n! has

OpenStudy (anonymous):

Go back to where we counted how many factors of p n! has

OpenStudy (anonymous):

Do you mean Like this: \[\frac{80}{3} + \frac{80}{9} + \frac{80}{27} = 26 + 8 + 2 = 36 ??\]

OpenStudy (anonymous):

We where each time we take the largest integer no greater then 80/3^j

OpenStudy (anonymous):

Yes, thats it

OpenStudy (anonymous):

Thanks buddy.. This is quite useful for me..

OpenStudy (anonymous):

ostow lawago...

OpenStudy (anonymous):

What does it mean @Tease //